OpenStudy (anonymous):

how do u expand and simplify this compound angle formula: cosec (A - 45 degrees)

OpenStudy (mathmale):

There are formulas for the sine, cosine and tangent of sums or differences of angles. You could also derive such a formula for the cosecant function, if you remember that csc x = 1 / (sin x). Thus, csc (a + b) = 1 / sin (a + b). Look up "sum and difference formulas for the cosine" for more details.